A random sample of 14 cigarettes of a certain brand has an average nicotine content of 3.7 milligrams and a standard deviation of 1.5 milligrams. What is the standard error of the 99% confidence interval for the average nicotine content of the cigarettes.



1
Expert's answer
2022-02-24T12:39:56-0500

The critical value for "\alpha = 0.01" and "df = n-1 = 13" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =3.012275."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

The standard error is


"SE_{\bar{X}} =\dfrac{s}{\sqrt{n}}=\dfrac{1.5}{\sqrt{14}}\approx0.4009"

The standard error of the 99% confidence interval for the average nicotine content of the cigarettes is 0.4009 milligrams.